OpenStudy (anonymous):

Given the functions f(n) = 11 and g(n) = -2(n - 1), combine them to create an arithmetic sequence, an, and solve for the 31st term. (2 points)

OpenStudy (anonymous):

an = 11 - 2(n - 1); a31 = -49 an = 11 - 2(n - 1); a31 = -51 an = 11 + 2(n - 1); a31 = 71 an = 11 + 2(n - 1); a31 = 73

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nincompoop

OpenStudy (haseeb96):

what do you think which one is correct ? @xangelaax

OpenStudy (anonymous):

1 or 2 but idk how to find it out

OpenStudy (haseeb96):

when we combine these functions we get 11 -2(n-1) ..... eq#1 now to find a31 put n=31 into eq#1 11-2(31-1) 11-2(30) 11-60 -49 is correct answer So 1 is correct answer

OpenStudy (anonymous):

thanks so much can you please help with 1 more?

OpenStudy (haseeb96):

yes but post it in the new question and click on best response if you like my explanation

OpenStudy (anonymous):

okay lol

OpenStudy (jiteshmeghwal9):

\[a_n=f(n)+g(n)\]\[a_n=11-2(n-1)\]\[a_{31}=11-2(31-1)\]

OpenStudy (anonymous):

thanks that makes sense

OpenStudy (anonymous):

i didnt mean to resend that ^^